# Challenging VLMs’ Structural Spatial Intelligence through Complex Reasoning Tasks

### v2

@@ 5 Conclusion
This paper presents SIRI-Bench, a benchmark designed to assess the structural spatial intelligence of VLMs through 3D mathematical reasoning problems. SIRI-Bench consists of 9078 video-question-answer triplets, where each question is formulated within a realistic 3D scene captured in video format. Solving these questions requires models to extract, interpret, and perform reasoning based on spatial cues, offering a novel and rigorous evaluation of VLMs’ spatial grounded reasoning ability. To construct this dataset efficiently, we develop an Automatic Scene Creation Engine that converts abstract math problems into realistic 3D scenes and renders corresponding videos. The engine coordinates multiple specialized LLM agents within a structured workflow, ensuring reliable scene generation while significantly reducing the cost of large-scale data production. Experiments on SIRI-Bench demonstrate that existing VLMs consistently fail to solve more than 50% problems when spatial information must be inferred from video. Their performance not only falls far behind human level but also lags significantly behind their text-only counterparts, which are provided with explicit mathematical conditions in text form. Further experiments reveal that existing VLMs over-rely on textual cues and struggle to actively and accurately extract key spatial information—such as object types, geometric dimensions, and spatial relations—from visual input. These findings highlight a fundamental limitation of current VLMs and demonstrate the value of SIRI-Bench***Upon acceptance, we will release the code and dataset for public access.
Ethics statement
Our work is conducted in accordance with ethical research practices and legal requirements. All original mathematical problems used in SIRI-Bench are licensed for research use, and the 3D scenes are generated based on the publicly available 3D-FRONT dataset, for which we have obtained proper authorization. The data generation process involves only synthetic scenes and automatically rendered videos, without collecting or inferring any personally identifiable information (PII). We emphasize that SIRI-Bench is designed solely for academic research on spatial reasoning and model evaluation. It should not be misused for surveillance, discriminatory applications, or other purposes that could harm individuals or communities.
Reproducibility Statement
We provide comprehensive details of our dataset construction pipeline. All prompts used in data generation and experiments are included in the Appendix to enable precise replication. We also describe our experimental settings, evaluation protocols, and ablation study configurations in detail. We will publicly release the SIRI-Bench dataset, along with the associated code and generation scripts, upon acceptance. This ensures that all results presented in the paper can be independently verified and extended by the research community.
LLM Clarification
All uses of Large Language Models (LLMs) in this work are explicitly documented in the paper. During dataset construction, they were used solely as auxiliary tools for repetitive tasks such as condition text editing and dimension solving, always under human supervision. Crucially, all core scientific contributions—including the formulation of the research problem, the design of the SIRI-Bench dataset, the development of the scene generation pipeline, and the experimental design and analysis—were conceived and executed by the authors. The use of LLMs did not involve conceptualization, methodology design, or result interpretation, and did not reach a level that would warrant considering them as contributors.

### v3

@@ 5 Conclusion
This paper presents SIRI-Bench, a benchmark designed to assess the structural spatial intelligence of VLMs through 3D mathematical reasoning problems. SIRI-Bench consists of 9078 video-question-answer triplets, where each question is formulated within a realistic 3D scene captured in video format. Solving these questions requires models to extract, interpret, and perform reasoning based on spatial cues, offering a novel and rigorous evaluation of VLMs’ spatial grounded reasoning ability. To construct this dataset efficiently, we develop an Automatic Scene Creation Engine that converts abstract math problems into realistic 3D scenes and renders corresponding videos. The engine coordinates multiple specialized LLM agents within a structured workflow, ensuring reliable scene generation while significantly reducing the cost of large-scale data production. Experiments on SIRI-Bench demonstrate that existing VLMs consistently fail to solve more than 50% problems when spatial information must be inferred from video. Their performance not only falls far behind human level but also lags significantly behind their text-only counterparts, which are provided with explicit mathematical conditions in text form. Further experiments reveal that existing VLMs over-rely on textual cues and struggle to actively and accurately extract key spatial information—such as object types, geometric dimensions, and spatial relations—from visual input. These findings highlight a fundamental limitation of current VLMs and demonstrate the value of SIRI-Bench.
Appendix A Data Reliability
Ensuring the reliability of data is essential for a benchmark designed to evaluate spatial reasoning. We validate the reliability of SIRI-Bench from three key perspectives: geometric dimension solving, condition text processing, and numerical answer correctness.
A.1 Geometric Dimension Solving
To assess the reliability of geometric dimensions solved by our Automatic Scene Creation Engine, we sample 70 cases (specifically, first 70 cases) from the dataset and asked human participants to evaluate the results. Each participant watch the rendered video of a scene and compared it against the original schematic diagram of the problem. A case is marked as correct if the rendered geometry matched the schematic and incorrect otherwise. The average correct rate across all cases is $81.4\%\pm 1.4\%$. Yet, the actual reliability is likely higher, since the schematic serves only as an illustrative reference and is often more restrictive than the original problem specification.
A.2 Condition Text Processing
To remove key spatial information from problem conditions while preserving auxiliary details, we employ Qwen-LM (Yang et al., 2024a) with a carefully designed prompt (see appendix F). The model is strictly instructed to remove only the information related to the primary geometric object, ensuring that essential supporting information remains intact. This design means that even if Qwen-LM occasionally leaves some targeted information undeleted, it will never mistakenly remove essential information, thus preserving the solvability of every problem.
Additionally, we replace letter-based vertex indices with color-based indices using a deterministic regular expression–based algorithm. The deterministic nature of this replacement process ensures the consistency and reliability of indices replacing.
A.3 Numerical Answer Correctness
Validating the correctness of mathematical answers is both critical and challenging. Existing tools such as math_verify huggingface (2025) can determine whether two symbolic expressions are mathematically equivalent. However, we choose to convert symbolic expressions into numerical values using Qwen-LM (Yang et al., 2024a). This decision is motivated by two considerations: (1) Previous work (including math_verify) reports comparable reliability to LLM-based approaches like Qwen-LM huggingface (2025). (2) Compared with true-or-false judgement, numerical outputs enable direct computation of relative errors, which is crucial for evaluating models’ spatial perception accuracy in our experiments.
Figure 11: Data Distribution. The SIRI-Bench dataset ensures a balanced distribution across geometric types (left), and problem types (right).
Appendix B Data Distribution
Figure 11 illustrates the distribution of problems within our dataset across different geometric types and problem types. The geometric entity types refer to the primary geometric entity involved in the problems, with a total of 13 distinct types. Problem types denote the specific quantities that the problems require solving for, including angle, ratio, length, area, volume. As depicted in fig. 11, our dataset achieves a balanced distribution across all geometric entity types and problem types. This balanced distribution ensures that our dataset can comprehensively evaluate the performance of VLMs on a wide range of problems, providing a robust benchmark for assessing their capabilities in various contexts.
Appendix C More Data Visualizations
We provide 70 additional data samples in the supplementary file, together with their corresponding videos and intermediate results. As can be consistently observed across these samples, our data generation engine accurately solves for geometric conditions, replaces vertex indices, processes textual conditions, and computes numerical answers, demonstrating its robustness and reliability.
Appendix D Why QVQ Performs So Poorly
As illustrated in Figure 1 of our main paper, QVQ Team (2024b) demonstrates the lowest performance among all evaluated visual-language models (VLMs) on the SIRI-Bench benchmark. Given QVQ’s design as a multimodal reasoning model with advanced visual understanding capabilities, this underperformance warrants further investigation.
QVQ Team (2024b) is built upon the Qwen2-VL-72BWang et al. (2024b) architecture and is tailored for complex visual reasoning tasks, including mathematical problem-solving. It has achieved commendable results on benchmarks such as MMMU, MathVista, and MathVision, showcasing its potential in handling intricate visual and mathematical challenges. Notably, QVQ is designed to emulate human-like reasoning by engaging in step-by-step analysis and self-reflection during problem-solving. However, when applied to SIRI-Bench, which presents novel 3D geometry problems requiring spatial reasoning, QVQ struggles significantly.
Upon examining its reasoning processes, we observe that QVQ often engages in extended chains of thought, spanning up to 8K or even 16K tokens, without arriving at a definitive conclusion. This behavior suggests that QVQ becomes entangled in recursive reasoning loops, failing to produce final answers. Consequently, its responses are categorized as having errors exceeding 200%, contributing to its poor performance metrics.
Appendix E Limitation and Future Work
Although our benchmark has uncovered several insightful findings under the current experimental settings, all videos are currently generated under a fixed camera configuration and consistent environment conditions. This design ensures a unified comparison but limits the diversity of visual input.
Given the flexibility of our data generation engine, it can easily create videos with varied object scales, backgrounds, and lighting conditions for the same underlying problem. Future work could introduce more diverse realistic 3D scenes to enable broader evaluations of VLMs It could also explore interactive settings where VLMs navigate 3D environments to acquire spatial information, which would be highly intriguing.
Appendix F Full Prompts
This section presents all the detailed prompts used in our study.
F.1 Inference Prompt
During VLM’s inference, we use the following prompt to guide model outputs:
Solve the following solid geometry problem. Think step by step, estimate and find the numerical answer. Use meters as the basic unit for length, and use radians for angles. In the last line of your answer, provide the final answer. The final answer should be a numerical value and must not contain formulas, symbols, unknown quantities, or references to unknown quantities. The last line of your response should be of the form Answer: \$ANSWER (without quotes) where \$ANSWER is the answer to the problem with no
boxed command.
{question}
Remember to put your answer on its own line after "Answer:", and do not use
boxed command.
F.2 Neural Translation Prompt
We use the following prompt to translate the mathematical problem from Chinese to English:
I will provide you with a math problem described in Chinese, usually a geometry problem. Your task is to translate this problem into English. You do not need to solve this math problem. You only need to translate it into English. Respond with the English translation directly, with no additional content.
Now, here is the Chinese geometry problem:
{question_zh}
Please translate it.
F.3 Dimension-Solving Prompt
In our Automatic Scene Creation Engine, we employ a math-specialist agent (Qwen-Math (Yang et al., 2024b) to solve for all key conditions necessary to fully define the main geometric entity.
I will provide you with a 3D geometric solid along with its conditions. Your task is to calculate the specific dimensions of this geometric solid. If the given conditions do not uniquely determine the solid, you may output any one valid solution that satisfies the constraints. Do not include units in your answer.
Here, the given 3D geometric solid is: {geometry_type}.
The given conditions of this geometric solid are:
{condition}.
Your task is to {solving_target}
After completing the reasoning and calculations, summarize the final result in the following format:
Final Answer: xxx
In your final answer, when giving values, please clearly label what they represent. For example:
Final Answer: ’The coordinate of [point name] is [value]’, ’The value of [dimension] is [value]’.
F.4 Condition-Processing Prompt
In our Automatic Scene Creation Engine, we refine the problem descriptions to retain only essential information while minimizing textual mathematics. This is achieved by an LLM agent (Qwen-LM (Yang et al., 2024a)), using the following prompt.
Please do the step by step reasoning.
You will be given a description that is part of a 3D geometry problem. In this description, there is a {geometry_type} that serves as the main geometric object (it may not always be referred to as ’{geometry_type}’ but could appear under other related names). And the description also contains other entities such as points or line segments.
Your task is to carefully determine which parts of the description describe the main geometric object and which describe other entities, then remove all information about the main geometric object while strictly preserving all information about the other entities.
Specifically, the main geometric object (i.e. the {geometry_type}) is defined by {n_vertex} vertices. Any points, lines, or planes composed exclusively of these {n_vertex} vertices belong to the main geometric object, and statements concerning only these elements must be removed, including their dimensions, shapes, geometric characteristics, etc. Any entities composed entirely of other elements, or partly of these {n_vertex} vertices together with other elements, are considered outside the main geometric object, and statements about them must be kept. If a statement involves both the main geometric object and other entities, keep it, but replace any names or references to the main geometric object with ’the object,’ ensuring the statement contains no words that reveal its shape (e.g., polyhedron, cube, prism, pyramid, frustum, etc.).
Please do the step by step reasoning; first correctly identify the main geometric object and the vertices that define it; then determine whether each entity belongs to the main geometric object or not; then split the passage into as fine-grained sub-statements as possible and carefully evaluate each sub-statement; then perform the removals and retentions; and finally summarize the processed result in the following format:
Final Answer: the description text with some information removed. (Do not include any explanations or enumeration symbols in the final answer, but rather provide a coherent descriptive sentence.)
Now, here is the given description:
{condition}.
Please process it.
F.5 Solving Numerical Answer Prompt
In order to measure prediction errors and analyze their distribution, we convert the answer from symbolic expressions to Python expressions by Qwen-LM (Yang et al., 2024a), using the following prompt. The Python expressions are then converted to numerical value.
You will be given a mathematical expression that may include formulas or symbols written in Markdown, LaTeX, or other formats. Your task is to accurately parse and convert the expression into a single valid Python expression using only the standard math library. Please do step-by-step reasoning and accurately carry out the conversion.
The following specific requirements apply:
- Do not include any units in the final expression.
- Treat angles as radians by default; if an angle is explicitly given in degrees, wrap it with math.radians(...).
- Do not use libraries other than math.
- After your reasoning, provide the final answer in this format, containing only the Python expression with no extra words, symbols, or units:
Final Answer: python_expression.
Now here is the given expression:
{answer}.
Please convert it.